how do you scale a triangle to 1/4 its size
Answer by TimothyLamb(4379)

---
scale each of its side by 1/4 like so:
---
new a = a(1/4)
new b = b(1/4)
new b = b(1/4)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php